OpenStudy (anonymous):

need help 4/9. - =8/27

OpenStudy (anonymous):

Is it \(\huge{\frac{4}{9}-\frac{8}{27}}\)

OpenStudy (anonymous):

Find the LCM of 9 & 27 For that Write down the Factors of both denominators (Number which is below the fraction) \(9 = 3 \times 3\)\[27= 3 \times 3 \times 3\] Now we do our largest count 3 x 3 x 3 =27 so 27 is the LCM Now we try to make our denominators equal (we can only subtract or add two fractions when their denominators are equal ) \[{{4 \times 3} \over { 9 \times 3 }}-{8 \over 27}= {12 \over 27}-{8 \over 27}={ {12-8}\over {27}}= {4 \over 27} \]